OpenStudy (anonymous):

what is the equation of the line that passes through (5, -2) and (-3, 4)

OpenStudy (anonymous):

Find the rate of change (slope). \[\frac{ 4-(-2) }{ -3-5 }=\frac{ 6 }{ -8} or \frac{ 3 }{ -4 }\] I am assuming you're using slope intercept form here. y=mx+b m=3/-4 Substitute a coordinate for x and y to solve for b. -2=3/-4(5)+b -2=15/-4+b Add 15/4 to both sides to solve for b. 15/4-2=b Find an LCD and simplify 15/4-8/4=b b=7/4 Your equation is y=3/-4x+7/4